# Robotic Learning for Informative Path Planning

### v1

@@ 5 Applications
Indoor exploration, point-goal navigation
UGV/UAV
Depth sensor
2D
Mapping, sensing
SL
(Ramakrishnan et al., 2020)
Indoor exploration
UGV/UAV
RGB-D camera
2D
Mapping, sensing
SL
(Gazani et al., 2023)
Infrastructure inspection
UAV
RGB-D camera
3D
RL
(Kumar et al., 2022)
Interactive exploration
Robot arm
RGB-D camera
3D
RL
(Chen et al., 2022)
Multi-object navigation
UGV
RGB-D camera
2D
Revolute joint camera
RL
(Novkovic et al., 2020)
Object finding
Robot arm
RGB-D camera
3D
Mapping, sensing
RL
(Best et al., 2019)
Object recognition, team orienteering
UGV team
2D Laser
2D
Mapping, sensing
RL
(Mendoza et al., 2020)
Object reconstruction
Any platform
Depth sensor
3D
SL
(Dhami et al., 2023)
Object reconstruction
Any platform
Depth sensor
3D
SL
(Vasquez-Gomez et al., 2021)
Object reconstruction
Any platform
Depth sensor
3D
Mapping
SL
(Pan et al., 2023b)
Object reconstruction
Any platform
RGB camera
3D
SL
(Pan et al., 2022)
Object reconstruction
Any platform
RGB camera
3D
Sensing
SL, AL
(Zhan et al., 2023)
Object reconstruction
Any platform
RGB camera
3D
Sensing
SL, AL
(Sünderhauf et al., 2023)
Object reconstruction
Any platform
RGB camera
3D
Sensing
SL, AL
(Jin et al., 2023)
Object reconstruction
Any platform
RGB camera
3D
Sensing
SL, AL
(Sukkar et al., 2019)
Object reconstruction
Robot arm
RGB-D camera
3D
Mapping, sensing
RL
(Lee et al., 2022)
Object reconstruction
Any platform
RGB-D camera
3D
Sensing
SL, AL
(Ran et al., 2022)
Object reconstruction
Any platform
RGB-D camera
3D
Sensing
SL, AL
(Yan et al., 2023)
Object reconstruction
Any platform
RGB-D camera
3D
Sensing
SL, AL
(Dhami et al., 2023)
Object reconstruction
UAV
Depth sensor
3D
ROS $*{}^{*}$
SL
(Pan et al., 2023a)
Object surface reconstruction
Any platform
RGB-D camera
3D
Sensing
ROS $*{}^{*}$
SL
(Cao et al., 2023c)
Persistent target monitoring
Robot team (UAV/UGV)
Generic camera
2D
Target position
SL, RL
(Ott et al., 2023)
Rover exploration
UGV
Drill, mass spectrometer
2D
Mapping, sensing
RL
(Choudhury et al., 2020)
Search and rescue
Any platform
Categorical survivor detection
2D
Mapping, sensing
RL
(Singh et al., 2009)
Search and rescue
Any platform
Optical sensor
2D
Sensing
RL
(Lim et al., 2016)
Search tasks, e.g., grasping pose or target search
Any platform
Binary measurements
3D
Sensing
(Reinhart et al., 2020)
Subterranean exploration
UAV
Depth sensor
3D
IL
(Qie et al., 2019)
Target assignment
UAV team
Not specified
2D
RL
(Gao and Zhang, 2022)
Target assignment
UAV team
Not specified
2D
Mapping, sensing
RL
(Hüttenrauch et al., 2019)
Target tracking
UAV team
Laser
2D
Mapping
Unicycle model
RL
(Ye et al., 2018)
Targeted object search
UGV
RGB camera
2D
Object recognition
RL
(Saroya et al., 2020)
Tunnel exploration
UGV
Depth sensor
2D
SL
(Ly and Tsai, 2019)
Urban exploration, surveillance
UGV/UAV
Depth sensor
2D
SL
(Hepp et al., 2018)
Urban exploration
UAV
Depth sensor
3D
Mapping, sensing
SL
(Zhang et al., 2023)
View planning for grasping
Any platform
RGB-D camera
3D
SL, AL
@@
Semantics
(Blum et al., 2019)
Land cover mapping
UAV
RGB camera
2D
Sensing
AL
(Georgakis et al., 2022b)
Indoor goal navigation
UGV
RGB-D camera
2D
Mapping, sensing
AL
(Zurbrügg et al., 2022)
Indoor semantic mapping
UGV
Laser, RGB-D camera
3D
Mapping, sensing, localization
AL
(Rückin et al., 2023)
Industrial and urban mapping, land cover analysis
UAV
RGB camera
2D
Mapping, sensing
(Rückin et al., 2022b)
Urban mapping
UAV
RGB camera
2D
Mapping, sensing
AL
(Rückin et al., 2024)
Urban mapping
UAV
RGB camera
2D
Mapping, sensing
AL
Active SLAM
(Hanlon et al., 2023)
Active visual localization
UGV
RGB-D camera
3D
Localization
SL
(Popović et al., 2020)
Indoor temperature monitoring
UGV/UAV
Laser, temperature sensor
2D/3D
Mapping, sensing, localization
SL
(Chen et al., 2020)
Exploration
UGV
Lidar
2D
Mapping, sensing, localization
RL
(Chen et al., 2021)
Exploration
UGV
Lidar
2D
Mapping, sensing, localization
RL
(Bartolomei et al., 2021)
Goal-based navigation
UAV
RGB-D camera
3D
Mapping, sensing, localization
Spline path
RL
@@ 6.1 Generalizability
Most of the AIPP methods surveyed were tested in the same or similar environments as those seen during training. The generalizability of deep learning-based systems is especially important because the neural networks that are often used in AIPP frameworks (Choudhury et al., 2018; Shrestha et al., 2019; Tao et al., 2023; Schmid et al., 2023; Li et al., 2020; Dai et al., 2022) are known to produce overconfident wrong predictions in out-of-distribution scenarios (Abdar et al., 2021). Future work should address these inherent generalization limitations to improve robustness. We identified one possible way to deal with deep learning-based vision uncertainties in Sec. 4.4 by combining uncertainty-aware deep learning and active learning approaches with AIPP (Chaplot et al., 2021; Zurbrügg et al., 2022; Rückin et al., 2023). Similarly, RL-based AIPP systems are increasingly popular (Cao et al., 2023a; Yang et al., 2023a; Cao et al., 2023b; Chen et al., 2021; Wei and Zheng, 2020; Choi and Cielniak, 2021). However, RL approaches degrade in performance when deployed in unseen simulated environments or real-world scenarios (Bousmalis et al., 2018; Zhao et al., 2020; Zhu et al., 2020). Future work on RL-based AIPP methods should focus on evaluating the learned policy in more diverse environments and in the real world. Further, new methods are needed to improve the generalization of RL for AIPP (Kirk et al., 2023). Other subfields of RL, such as meta-reinforcement learning (Gupta et al., 2018), could serve as inspiration for enabling efficient online policy adaption to unseen environments. Data augmentation and domain randomization during training could improve test-time policy robustness (Sadeghi and Levine, 2017).
@@ 6.2 Handling Localization Uncertainty
Effectively handling robot localization uncertainty is a key aspect in advancing learning-based AIPP in robotics. Localization uncertainties arise from various sources, including sensor noise, environmental changes, and the inherent limitations of localization technologies. However, most existing approaches assume perfect knowledge of the robot pose, both during planning and plan execution, which limits their practical applicability. Some works in active SLAM account for localization uncertainty via belief space planning (Chen et al., 2020, 2021) or modifying the planning objective (Popović et al., 2020) using graph-based representations of the robot state. Incorporating such probabilistic methods into learning-based AIPP methods and exploring transfer learning techniques (Chen et al., 2021; Genc et al., 2020) can enhance robustness.
@@ 6.3 Temporal Changes and Long-Term Planning
Addressing temporal changes and ensuring robust long-term planning in learning-based AIPP poses a critical challenge. Temporal changes, such as variations in environmental conditions, seasonal shifts, or the presence of dynamic obstacles, require AIPP methods to not only adapt to the immediate surroundings but also plan for sustained effectiveness over long durations. This involves developing strategies that consider the evolution of the environment and adjust the planned paths online. Whereas several works study spatiotemporal monitoring (Caley et al., 2019; Marchant et al., 2014; Cao et al., 2023c) and dynamic target tracking (Tzes et al., 2023; Hüttenrauch et al., 2019), there is a need for new approaches that can handle multiple moving objects in 3D environments. We also highlight the importance of developing such methods for agile platforms such as small UAV platforms while also accounting for their physical motion constraints, given that they often operate in dynamic and changing landscapes.
@@ 6.4 Heterogeneity
One challenge is the development of learning-based AIPP methods that exploit the benefits of heterogeneous robot teams or heterogeneous sensor modalities within a single robot. Leveraging heterogeneity in robotic systems has a high potential to improve sensing and monitoring performance. For instance, a tandem robotic system can combine the agility of a small UAV and the high-payload, detailed inspection capabilities of a UGV. A robot equipped with different sensors, e.g., a drill and a mass spectrometer (Ott et al., 2023), can measure different physical parameters during a single mission. To fully exploit these capabilities, we recommend investigating principled ways of incorporating heterogeneity, also considering collaboration aspects for effective decision-making.
@@ 7 Conclusion
In this survey paper, we analyzed the current research on AIPP in robotic applications, focusing especially on methods that exploit recent advances in robot learning. We introduced a unified mathematical framework for addressing problems in learning-based AIPP. Based on this formulation, we provided two complementary taxonomies from the perspectives of (i) learning algorithms and (ii) robotic applications. We provided a structured outline of the research landscape and highlighted synergies and trends. Finally, we discussed the remaining challenges for learning-based AIPP in robotics and outlined promising research directions.

### v3

@@ 5 Applications
[36] Indoor exploration, point-goal navigation UGV/UAV Depth sensor 2D Mapping, sensing - SL
[37] Indoor exploration UGV/UAV RGB-D camera 2D Mapping, sensing - SL
[93] Infrastructure inspection UAV RGB-D camera 3D - - RL
[98] Interactive exploration Robot arm RGB-D camera 3D - - RL
[50] Multi-object navigation UGV RGB-D camera 2D - Revolute joint camera RL
[139] Object finding Robot arm RGB-D camera 3D Mapping, sensing - RL
[75] Object recognition, team orienteering UGV team 2D Laser 2D Mapping, sensing - RL
[48] Object reconstruction Any platform Depth sensor 3D - - SL
[57] Object reconstruction Any platform Depth sensor 3D - - SL
[47] Object reconstruction Any platform Depth sensor 3D Mapping - SL
[82] Object reconstruction Any platform RGB camera 3D - - SL
[78] Object reconstruction Any platform RGB camera 3D Sensing - SL, AL
[80] Object reconstruction Any platform RGB camera 3D Sensing - SL, AL
[81] Object reconstruction Any platform RGB camera 3D Sensing - SL, AL
[77] Object reconstruction Any platform RGB camera 3D Sensing - SL, AL
[30] Object reconstruction Robot arm RGB-D camera 3D Mapping, sensing - RL
[84] Object reconstruction Any platform RGB-D camera 3D Sensing - SL, AL
[79] Object reconstruction Any platform RGB-D camera 3D Sensing - SL, AL
[83] Object reconstruction Any platform RGB-D camera 3D Sensing - SL, AL
[57] Object reconstruction UAV Depth sensor 3D - ROS MoveIt SL
[58] Object surface reconstruction Any platform RGB-D camera 3D Sensing ROS MoveIt SL
[70] Persistent target monitoring Robot team (UAV/UGV) Generic camera 2D Target position - SL, RL
[62] Rover exploration UGV Drill, mass spectrometer 2D Mapping, sensing - RL
[7] Search and rescue Any platform Categorical survivor detection 2D Mapping, sensing - RL
[8] Search and rescue Any platform Optical sensor 2D Sensing - RL
[104] Search tasks, e.g., grasping pose or target search Any platform Binary measurements 3D Sensing - -
[42] Subterranean exploration UAV Depth sensor 3D - - IL
[124] Target assignment UAV team Not specified 2D - - RL
[53] Target assignment UAV team Not specified 2D Mapping, sensing - RL
[72] Target tracking UAV team Laser 2D Mapping Unicycle model RL
[130] Targeted object search UGV RGB camera 2D Object recognition - RL
[96] Tunnel exploration UGV Depth sensor 2D - - SL
[107] Urban exploration, surveillance UGV/UAV Depth sensor 2D - - SL
[56] Urban exploration UAV Depth sensor 3D Mapping, sensing - SL
[97] View planning for grasping Any platform RGB-D camera 3D - - SL, AL
@@
Semantics [88] Land cover mapping UAV RGB camera 2D Sensing - AL
[94] Indoor goal navigation UGV RGB-D camera 2D Mapping, sensing - AL
[89] Indoor semantic mapping UGV Laser, RGB-D camera 3D Mapping, sensing, localization - AL
[90] Industrial and urban mapping, land cover analysis UAV RGB camera 2D Mapping, sensing -
[134] Urban mapping UAV RGB camera 2D Mapping, sensing - AL
[91] Urban mapping UAV RGB camera 2D Mapping, sensing - AL
Active SLAM [108] Active visual localization UGV RGB-D camera 3D Localization - SL
[61] Indoor temperature monitoring UGV/UAV Laser, temperature sensor 2D/3D Mapping, sensing, localization - SL
[29] Exploration UGV Lidar 2D Mapping, sensing, localization - RL
[34] Exploration UGV Lidar 2D Mapping, sensing, localization - RL
[122] Goal-based navigation UAV RGB-D camera 3D Mapping, sensing, localization Spline path RL
@@ 6.1 Generalizability
Most of the AIPP methods surveyed were tested in the same or similar environments as those seen during training. The generalizability of deep learning-based systems is especially important because the neural networks that are often used in AIPP frameworks [31, 44, 51, 38, 106, 55] are known to produce overconfident wrong predictions in out-of-distribution scenarios [135]. Future work should address these inherent generalization limitations to improve robustness. We identified one possible way to deal with deep learning-based vision uncertainties in Sec. 4.4 by combining uncertainty-aware deep learning and active learning approaches with AIPP [92, 89, 90]. Similarly, RL-based AIPP systems are increasingly popular [2, 63, 41, 34, 66, 69]. However, RL approaches degrade in performance when deployed in unseen simulated environments or real-world scenarios [140, 141, 142]. Future work on RL-based AIPP methods should focus on evaluating the learned policy in more diverse environments and in the real world. Further, new methods are needed to improve the generalization of RL for AIPP [143]. Other subfields of RL, such as meta-reinforcement learning [144], could serve as inspiration for enabling efficient online policy adaption to unseen environments. Data augmentation and domain randomization during training could improve test-time policy robustness [145].
@@ 6.2 Handling Localization Uncertainty
Effectively handling robot localization uncertainty is a key aspect in advancing learning-based AIPP in robotics. Localization uncertainties arise from various sources, including sensor noise, environmental changes, and the inherent limitations of localization technologies. However, most existing approaches assume perfect knowledge of the robot pose, both during planning and plan execution, which limits their practical applicability. Some works in active SLAM account for localization uncertainty via belief space planning [29, 34] or modifying the planning objective [61] using graph-based representations of the robot state. Incorporating such probabilistic methods into learning-based AIPP methods and exploring transfer learning techniques [34, 146] can enhance robustness.
@@ 6.3 Temporal Changes and Long-Term Planning
Addressing temporal changes and ensuring robust long-term planning in learning-based AIPP poses a critical challenge. Temporal changes, such as variations in environmental conditions, seasonal shifts, or the presence of dynamic obstacles, require AIPP methods to not only adapt to the immediate surroundings but also plan for sustained effectiveness over long durations. This involves developing strategies that consider the evolution of the environment and adjust the planned paths online. Whereas several works study spatiotemporal monitoring [49, 71, 70] and dynamic target tracking [74, 72], there is a need for new approaches that can handle multiple moving objects in 3D environments. We also highlight the importance of developing such methods for agile platforms such as small UAV platforms while also accounting for their physical motion constraints, given that they often operate in dynamic and changing landscapes.
@@ 6.4 Heterogeneity
One challenge is the development of learning-based AIPP methods that exploit the benefits of heterogeneous robot teams or heterogeneous sensor modalities within a single robot. Leveraging heterogeneity in robotic systems has a high potential to improve sensing and monitoring performance. For instance, a tandem robotic system can combine the agility of a small UAV and the high-payload, detailed inspection capabilities of a UGV. A robot equipped with different sensors, e.g., a drill and a mass spectrometer [62], can measure different physical parameters during a single mission. To fully exploit these capabilities, we recommend investigating principled ways of incorporating heterogeneity, also considering collaboration aspects for effective decision-making.
@@ 7 Conclusion
In this survey paper, we analyzed the current research on AIPP in robotic applications, focusing especially on methods that exploit recent advances in robot learning. We introduced a unified mathematical problem definition for addressing tasks in learning-based AIPP. Based on this formulation, we provided two complementary taxonomies from the perspectives of (i) learning algorithms and (ii) robotic applications. We provided a structured outline of the research landscape and highlighted synergies and trends. Finally, we discussed the remaining challenges for learning-based AIPP in robotics and outlined promising research directions.
@@ 8 Funding Sources
We would like to thank Jonas Schaap for helping with the infographics design.